OpenStudy (anonymous):

simplify the expression. 2(6b+9m-8m)

mathslover (mathslover):

\[\huge{a(b+c+d)=ab+ac+ad}\] \[\huge{2(6b+9m-8m)=2(6b)+2(9m)-2(8m)}\] or : \[\huge{2(6b+9m-8m)=2(6b+m)=2(6b)+2(m)}\]

OpenStudy (maheshmeghwal9):

catch the like terms one side & unlike terms .............. omg @mathslover gt it lol :D